How many liters of 22% and 12% acetic acid should be mixed to produce 48 liters of 15% acetic acid?
Answer by Earlsdon(6294)

Let x = the number of liters of 22% acetic acid solution required, then (48-x) = the number of liters of 12% acetic acid solution required.
The sum of these two amounts = 48 liters of 15% acetic acid solution.
Putting this into algebra and converting the percentages to their decimal equivalents, we get:
0.22x + (0.12(48-x) = 0.15(48) Simplify and solve for x.
0.22x + 5.76 - 0.12x = 7.2
0.1x + 5.76 = 7.2 Subtract 5.76 from both sides.
0.1x = 1.44 Divide both sides by 0.1
x = 14.4 liters of 22% acetic acid solution.
48-x = 48-14.4 = 33.6 liters of 12% acetic acid solution.